OpenStudy (anonymous):

Simplify log5 s + log5 11 - log5 15 - log5 t

OpenStudy (anonymous):

Remember this rule:\[\log(a _{1})+\log(a _{2})+\log(a _{3})...+\log(a_{n})=\log(a _{1}*a _{2}*a _{3}...*a _{n})\] @peaceloveeat

OpenStudy (anonymous):

I haven't learned any of that yet, they want me to use the power property of logarithms, but I don't understand how to start it. @genius12

OpenStudy (anonymous):

The power property goes like this:\[\log(a^2)=2\log(a)\]It just means that if the thing you're taking the logarithm of has an exponent, we can bring that exponent to the front and multiply it by the log instead. @peaceloveeat

OpenStudy (anonymous):

An example of Power property:\[\log(10^2)=2\log(10)=2\]An example of sum property:\[\log(4)+\log(50)+\log(5)=\log(4*50*5)=\log(1000)=3\]Usually, when the base of the logarithm is not stated, it is assumed that the base is 10. Now can you apply these properties to solve the given problem? It's pretty simple. @peaceloveeat

OpenStudy (anonymous):

Ohh okay! I understand it now thanks! @genius12